OpenStudy (anonymous):

What is the end behavior of the function f(x) = x3 + 2x2 + 4x + 5? Up on the left, up on the right Up on the left, down on the right Down on the left, up on the right Down on the left, down on the right

OpenStudy (anonymous):

@willreel @quixoticideals @surjithayer @Awesome781 @Kenziehulse @Niall @xxccjxx @xxbatmanbetchesxx @kewlgeek555

OpenStudy (anonymous):

@willreel @Awesome781 @AravindG @kmacdaddy

OpenStudy (anonymous):

@willreel do you think you can help me

OpenStudy (anonymous):

@AravindG @FlvsGirl @Glizzy_Nedd

OpenStudy (anonymous):

so you think you can help me

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@wilstanc @wio @jhonyy9 @Jonathan1997 @LastDayWork @Niall

OpenStudy (anonymous):

I would help you, if you tell me what kind of math this is

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

@Jonathan1997

OpenStudy (anonymous):

Ok, well then i should know this cause i am in Algebra 2! :P lol

OpenStudy (anonymous):

lol so could you help

OpenStudy (anonymous):

https://s3.amazonaws.com/grapher/exports/uhhquxatcn.png Ok, so does this help you at all?

OpenStudy (anonymous):

And if not, then I can get you the answer in like 5 minuets, also, is the problem f(x) = x^3 + 2x^2 + 4x + 5? I added the carrot tops (^)

OpenStudy (anonymous):

no that didnt help but i believe its the last answer tho and ok i will be here waiting

OpenStudy (anonymous):

My Math teacher said it is C

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

@Jonathan1997 can you help me with one more question please

OpenStudy (anonymous):

Yeah, i can try

OpenStudy (anonymous):

What are the coordinates of the turning point for the function f(x) = (x + 2)3 - 4? (-2, -4) (-2, 4) (2, -4) (2, 4)

OpenStudy (anonymous):

this is the last question i promise

OpenStudy (anonymous):

https://www.desmos.com/calculator/7oiea4vkrz I get this... But the points are -0.4, and 4

OpenStudy (anonymous):

wait hoe did you get -0.4 thats not even an answer choice

OpenStudy (anonymous):

I know, that is what I don't get, i put it in the calculator like you said.... f(x) = (x+2)^3 - 4

OpenStudy (anonymous):

mmmm...do you think you could ask your math teacher

OpenStudy (anonymous):

Yes, i will right now

OpenStudy (anonymous):

thank you so much your such a lifesaver

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Jonathan1997 where did you go

OpenStudy (anonymous):

did your math teacher solve it yet

OpenStudy (anonymous):

@Jonathan1997

OpenStudy (anonymous):

Is it to the 3rd or second power

OpenStudy (anonymous):

are you talking about the problem or the answer

OpenStudy (anonymous):

The problem

OpenStudy (anonymous):

f(x) = x3 + 2x2 - 5x - 6

OpenStudy (anonymous):

No, the second one

OpenStudy (anonymous):

that is the second 1

OpenStudy (anonymous):

I am going to go with B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

r you sure

OpenStudy (anonymous):

Yeh, should be, If not then you can take my medal

OpenStudy (anonymous):

yea i had got the answer wrong

OpenStudy (anonymous):

but its ok i still got a good grade on it